OpenStudy (priyar):

A parallel plate capacitor has a capacitance C and charge Q.Another plate is inserted in between the plates of the capacitor at the center and is connected by a long wire to far off spherical conductor of radius R. What is the potential difference between the edge of the sphere and one of the outer plates?

OpenStudy (priyar):

@agent0smith @rvc

OpenStudy (priyar):

@ganeshie8 @hartnn

OpenStudy (priyar):

my answer is not matching..

OpenStudy (priyar):

I have a doubt.. why have they mentioned that the spherical capacitor is "far off" ? Is it of any significance..?

OpenStudy (priyar):

@Michele_Laino @Vincent-Lyon.Fr @IrishBoy123

OpenStudy (priyar):

@mayankdevnani

OpenStudy (priyar):

is it to specify that the spherical capacitor is isolated ?

OpenStudy (mayankdevnani):

is the answer Q/C ?

OpenStudy (priyar):

@mayankdevnani no..that's not the answer..

OpenStudy (mayankdevnani):

great !

OpenStudy (mayankdevnani):

let me try this again and get you back soon !

OpenStudy (priyar):

okay

OpenStudy (priyar):

@samigupta8

OpenStudy (anonymous):

V=CQ/R

OpenStudy (samigupta8):

Plate inserted between the parallel plates of capacitor .. Ryt?

OpenStudy (anonymous):

Since you assume the electric field to be homogeneous between the two fields, it comed out to just E = V / d, where V is the voltage between the two plates and d is the distance between them. This field is cause by some constant charge density ρ ( ρ = Q / A, where A is the area of the plate) - in this case E = ρ / ε. The voltage is the line integral of the electric field between one of the plates and the inserted one. \[V = \int\limits_0^{\frac{d}{2}} E \;dz\] Assuming your z-axis perpendicular to the plates and one of them is at z=0. This integral is completely trivial and you just need to plug in what we stated before.

OpenStudy (priyar):

yes @samigupta8 .. then?

OpenStudy (anonymous):

You're trying to find V, so substituting E for V / d will not get you anywhere, will it? :)

OpenStudy (priyar):

no.. i didn't mean like that.... i got it thanks !

OpenStudy (samigupta8):

Did u get it?? How??

OpenStudy (priyar):

Q/2C is the answer!

OpenStudy (priyar):

@MuH4hA 's method is correct.. simply put it is V= Q/ C' where C' is the new capacitance.. which is 2C since d has reduced by half.. so V= Q/2C.. that's it..!

OpenStudy (priyar):

got it? @samigupta8

OpenStudy (samigupta8):

Okk bt priyar v have to like calculate the potential diff between the outer plate and sphere .. So don't you think that capacitance will be like the series combination of three capacitor ..

OpenStudy (samigupta8):

Like first capacitor is in between first plate(lhs) and the other plate's lhs Then between the other plate's .. And the last one between the other plate's (rhs) and the first plate (rhs)

OpenStudy (anonymous):

Yes - you can just make a symmetry argument and be done with it. However, this only works for an ideal plate-capacitor. (large plate surface in relation to d, so we can ignore the inhomogeneities at the edges and a constant ε). Actually solving the integral would work for other geometries as well - but it might be quite hard.

OpenStudy (priyar):

i will tell you later

OpenStudy (samigupta8):

@muh4ha will u plss explain clearly?

OpenStudy (anonymous):

@samigupta8 Explain the symmetry argument? Or the integral? What seems to be the problem? :)

OpenStudy (samigupta8):

Well! I m not linking anything with your way... I just wanna make sure if we can solve for the three capacitance and solve fof the net capacitance and get the value of charge

OpenStudy (samigupta8):

For*

OpenStudy (anonymous):

I'm not sure I follow (what's the "three capacitance" ?).. are you trying to think of it as 2 capacitors in series? \[\frac{1}{C_{total}} = \frac{V}{Q} = \frac{V_1 + V_2}{Q} = \frac{V_1}{Q} + \frac{V_2}{Q} = \frac{1}{C_1} + \frac{1}{C_2}\] or more general for N capacitors in series: \[\frac{1}{C_{total}} = \sum_{n = 1}^N \frac{1}{C_n}\] (like N parallel resistors). If you think about two connected capacitors in series, in the middle between them will be two plates (one of each capacitor) connected by a conductor. These two plates are physically disconnected from the rest of your electric circuit and thus the charge on them must remain constant. If we assume there is 0 charge on them, when no voltage is applied over the whole of the 2 capacitors, there'll still be 0 charge when some voltage is there. Therefore, the two connected plates in the middle must have either 0, or equal and opposite charge +/-Q from each other (to cancel out). Then, if you look at the capacitors on their own, since one plate will have the charge +/- Q, the other plate must have the opposite charge and you end up with both capacitors carrying the same charge Q. The voltage drops are of course different in general, but the total drop will be the sum of the two (since they are in series) and we can write V/Q = (V1 + V2) / Q. In our example, the two capacitors are identical (plate is inserted exaclty in the middle), therefore they will have to have the same capacitance, same stored charge and there'll be the same voltage drop on the individual capacitors.

OpenStudy (michele_laino):

here is my reasoning: the situation, is like below: